OpenStudy (hanner_b_nanner):

Can someone PLEASE help me??? Fan&&Medal! :)) Find the non-extraneous solutions of

OpenStudy (hanner_b_nanner):

 A) x = −2 B) x = −6 C) x = −6 and x = −2 D) x = 2

OpenStudy (hanner_b_nanner):

where the question things are, it's supposed to be negative signs..

OpenStudy (princeharryyy):

sqrt(x+6) = x+6 => square both sides x+6 = x^2 + 36 + 12x x^2 + 11x + 30 = 0

OpenStudy (princeharryyy):

further solve the equation to get values of x if it satisfies it isn't the extraneous solution if it doesn't than it is

OpenStudy (hanner_b_nanner):

That's what I got, but I'm totally confused on how the answer will be -2, 2, or -6... :(

OpenStudy (princeharryyy):

-6 would b ur answer @hanner_B_nanner

OpenStudy (princeharryyy):

(x+5)(x+6) = 0 => x=-5 or x = -6 @hanner_B_nanner no other solution exists.